OpenStudy (anonymous):

completing the square of -x^2 - 2x +3

OpenStudy (anonymous):

I got -(x+1)^2-4 but apparently its +4

hartnn (hartnn):

how did you get -4?

hartnn (hartnn):

\(-(x^2 + 2x+1 -4 ) \\ = -([x+1]^2 - 4) \\ = -[x+1]^2 -(-4) \\ = -(x+1)^2 +4\)

OpenStudy (danjs):

(x + a)^2 = x^2 + 2a*x + a^2 notice the relationship from 2a and a^2, that is where you take half your middle coeffiecient 1/2*[2a] = a, then square it , a^2

OpenStudy (danjs):

-x^2 - 2x +3 -( x^2 + 2x ) - 3 -------------------- 2a = 2 here

OpenStudy (anonymous):